OpenStudy (anonymous):

A 0.02864 g sample of vitamin A, a compound consisting only of carbon, hydrogen, and oxygen, is burned to form 0.08802 g of carbon dioxide and 0.02703 g of water. Calculate the percentage of each element in vitamin A.

OpenStudy (anonymous):

For the carbon dioxide -- find its percent composition of carbon. \[\color{green}{\% \text{C} = \frac{12.0~\text{g}}{44.0~\text{g}} \times 100 \approx 27.3\% ~\text{C} }\] This means that 27.3% of the 0.088 02 g of carbon dioxide is carbon: \[\color{green}{27.3\% \times 0.088\,02~\text{g} = 0.024\,0~\text{g of carbon}}\] So, the percent carbon in the original vitamin A sample is: \[\color{green}{\frac{0.024\;0~\text{g of carbon} } {0.028\,64~\text{g of vitamin A} } \times 100 = 83.7\% ~\text{carbon}}\] Follow a similar procedure to find the percent of hydrogen in the original sample. Once you know the percent of carbon and hydrogen, you can also determine the precent of oxygen, since all three percentages have to add up to 100%.